OpenStudy (anonymous):

a = 40, c = 41, b = use pythagorean theorem

OpenStudy (anonymous):

9

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

The Pythagorean theorem states: \[a^2 + b^2 = c^2\] So plug our values in: \[40^2 + b^2 = 41^2\] So: \[b=\sqrt{41^2 - 40^2} = \sqrt{81} = 9\]